OpenStudy (anonymous):

If 4 is added to twice a number and this sum is multiplied by 5, the result is the same as if the number is multiplied by -8 and 2 is added to the product. What is the number?

OpenStudy (mokeira):

let the number be x 4 is added to twice a number= 4+2x the sum is multiplied by 5= 5(4+2x) the result is the same as if the number is multiplied by -8 and 2 added= 5(4+2x)= (-8x)+2 so now you have 5(4+2x)=(-8x)+2 so now solve for x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Is -1 the answer

OpenStudy (mokeira):

5(4+2x)=(-8x)+2 first open up the brackets 20+10x=-8x+2 collect like terms 10x+8x=2-20 now simplify

OpenStudy (mokeira):

oh sorry you are right

OpenStudy (smart8thgrader):

-1 is the answer